OpenStudy (anonymous):

A boat has 2 motors, one large and one small. If only the large motor is running, a tank of fuel lasts 2 h. If only the small motor is running, the tank of fuel lasts 4 h. How long will the fuel last if both motors are running?

OpenStudy (raden):

short cut ! t=(2*4)/(2+4)=....? t=time

OpenStudy (anonymous):

thank you ! is there any other way of solving this?

OpenStudy (raden):

welcome,, t1+t2=1/2 + 1/4 = 2/4+1/4 = 3/4 = 1/(4/3) so, both motor are running 4/3 hours = 4/3 * 60 minutes = 80 minutes = 1 hour + 20 m